### v7

@@ 4.1 Preliminary Lemmas
Applying Gronwall’s lemma to the above inequality, we infer that
$\max_{t\in[0,T]}\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})}+\left[\mathbb{E}\int_{0}^{T}\|\eta_{h}(t)\|_{L^{2}}^{p-2}\|\eta_{h}(t)\|_{\dot{H}_{h}^{1,2}}^{2}\,\mathrm{d}t\right]^{1/p}\leqslant ch^{\alpha/p}.$
Moreover, for any $\epsilon>0$ and $t\in(0,T]$, we establish the following bounds:
$\lVert{(z-P_{h}z)(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant ch^{1-\epsilon}$ (by Eqs. 4.19 and 3.4),
@@
Since the equality Eq. 4.43 implies
@@
$\qquad{}+\|\xi_{h}(t)\|_{L^{p}(\Omega;L^{2})}+\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})},\quad\forall t\in[0,T],$
the combination of these bounds yields the desired error estimate Eq. 3.6. This completes the proof of Theorem 3.2.
@@
Proof of Theorem 3.3. Let $G$ be defined by Eq. 2.10, and let $z$ be defined by Eq. 4.34. Extending the reasoning from the proof of Theorem 3.1, we establish the following enhanced regularity properties for the processes $y$ and $z$:
$y\in L^{p}(\Omega;C([0,T];\dot{H}^{2-\epsilon,q})),$ $\forall p\in(2,\infty),\,\forall q\in[2,\infty),\,\forall\epsilon>0;$ (4.45)
$z\in L^{p}(\Omega\times(0,T);\dot{H}^{2,q}),$ $\forall p\in(2,\infty),\,\forall q\in[2,\infty);$ (4.46)
$z\in L^{p}(\Omega;C([0,T];\dot{H}^{2-\epsilon,q})),$ $\forall p\in(2,\infty),\,\forall q\in[2,\infty),\,\forall\epsilon>0.$ (4.47)
Let $G_{h}$ be defined by Eq. 4.36. We adopt the same notation as introduced in the proof of Theorem 3.2:
@@
The remainder of the proof is structured into the following five steps.
@@
$\left\|z-P_{h}z\right\|_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty),$ (4.48)
$\left\|G-G_{h}\right\|_{C([0,T];L^{q})}+\left\|P_{h}G-G_{h}\right\|_{C([0,T];L^{q})}\leqslant ch^{2},\quad\forall q\in[2,\infty).$ (4.49)
Estimate (4.48) is a direct consequence of (4.19) combined with the regularity property (4.46). Estimate (4.49) follows from (4.22) and the condition $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$. In analogy to the estimate (4.29), we utilize Lemma 4.7, along with (4.20) and the regularity assertion (4.46), to derive the following bound for $\xi_{h}$:
$\left\|\xi_{h}\right\|_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty).$ (4.50)
Following the reasoning from Theorem 3.2, we obtain the following estimate for $\eta_{h}$:
$\max_{t\in[0,T]}\left\|\eta_{h}(t)\right\|_{L^{p}(\Omega;L^{2})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.51)
In light of the decomposition Eq. 4.43, combining Eqs. 4.50, 4.51, 4.49 and 4.48 leads to the error estimate
@@
Furthermore, for any $p\in(2,\infty)$ and $q\in[2,\infty)$, we deduce the following chain of inequalities:
$\lVert{y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$
@@
$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}{}$ $\lVert{P_{h}y}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}+ch^{3(1/q-1/2)}\lVert{P_{h}(y-y_{h})}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}{}$ $c\lVert{y}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}+ch^{3(1/q-1/2)}\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
$\stackrel{{\scriptstyle\mathrm{(iii)}}}{{\leqslant}}{}$ $c\lVert{y}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}+ch^{2+3(1/q-1/2)},$
where in (i) we apply the inverse estimate $\lVert{v_{h}}\rVert_{L^{q}}\leqslant ch^{3(1/q-1/2)}\lVert{v_{h}}\rVert_{L^{2}}$ for all $v_{h}\in X_{h}$; in (ii) we utilize the stability property of $P_{h}$ as stated in Eq. 4.18; and in (iii) we employ the error estimate Eq. 4.52. By invoking the regularity assertion Eq. 4.45, we subsequently conclude a uniform boundedness with respect to the spatial mesh size $h$:
$\lVert{y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\text{ is uniformly bounded with respect to $h$ for all $p,q\in[2,\infty)$}.$ (4.53)
Step 2. Let us prove
$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.54)
Using the equality Eq. 4.38, along with the definitions of the discrete convolution operators $S_{0,h}$ and $S_{1,h}$ given in Eqs. 4.24 and 4.25, we deduce the following decomposition:
$\eta_{h}=II_{1}+II_{2}+II_{3},$ (4.55)
@@
$II_{1}$ $:=S_{0,h}\big{(}\xi_{h}+\eta_{h}+G_{h}-P_{h}G\big{)},$
$II_{2}$ $:=S_{0,h}P_{h}(y^{3}-y_{h}^{3}),$
$II_{3}$ $:=S_{1,h}P_{h}\big{(}F(y_{h})-F(y)\big{)}.$
Fix an arbitrary $p\in(2,\infty)$. For the term $II_{1}$, we establish the following bound for any $q\in[2,\infty)$:
$\lVert{II_{1}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$ $\leqslant c\lVert{S_{0,h}\big{(}\xi_{h}+\eta_{h}+G_{h}-P_{h}G\big{)}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{2,2})}$
$\leqslant c\lVert{\xi_{h}+\eta_{h}+G_{h}-P_{h}G}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
$\leqslant ch^{2},$
where the first inequality uses the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{\infty}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{2,2}}$ for all $v_{h}\in X_{h}$; the subsequent inequality relies on Lemma 4.7; and the final inequality is justified by Eqs. 4.51, 4.49 and 4.50. Turning to $II_{2}$, we establish the bound: for any $q\in[2,\infty)$,
$\lVert{II_{2}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$ $\leqslant c\lVert{S_{0,h}P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{2,5/3})}$
$\leqslant c\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{5/3})}$
$\leqslant c\lVert{y-y_{h}}\rVert_{L^{2p}(\Omega\times(0,T);L^{2})}\Big{(}\lVert{y}\rVert_{L^{4p}(\Omega\times(0,T);L^{20})}^{2}+\lVert{y_{h}}\rVert_{L^{4p}(\Omega\times(0,T);L^{20})}^{2}\Big{)}$
$\leqslant ch^{2},$
where the first inequality uses the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{\infty}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{2,5/3}}$ for all $v_{h}\in X_{h}$; the second inequality uses Lemma 4.7 and Eq. 4.18; the third inequality uses Hölder’s inequality; and the last inequality uses Eqs. 4.45, 4.52 and 4.53. For $II_{3}$ we establish the following bound:
$\lVert{II_{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$ $\leqslant c\lVert{S_{1,h}P_{h}\big{(}F(y_{h})-F(y)\big{)}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{1,2})}$
$\leqslant c\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
[…]
where the first inequality employs the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{6}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{1,2}}$ for all $v_{h}\in X_{h}$, and the second inequality invokes Lemma 4.8, along with Eqs. 4.18 and 4.4. In view of the decomposition Eq. 4.55, combining the above bounds for $II_{1}$, $II_{2}$, and $II_{3}$ yields Eq. 4.54.
Step 3. We proceed to establish the error estimate (3.7). Considering the decomposition of $y_{h}-y$ given in (4.43), we integrate the inequalities (4.48), (4.49), (4.50), and (4.54) to validate (3.7) for the specific case where $p\in(2,\infty)$ and $q=6$. Building upon this outcome and employing a similar bounding technique as utilized for $II_{3}$ in Step 2, we obtain an enhanced bound for $II_{3}$:
@@
$\lVert{II_{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty).$
Integrating this refined bound with the previously established bounds for $II_{1}$ and $II_{2}$ from Step 2, and incorporating the decomposition (4.55), we arrive at the following conclusion:
$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty).$ (4.56)
Ultimately, revisiting the decomposition (4.43) and synthesizing the inequalities (4.48), (4.49), (4.50), and (4.56), we affirm the validity of the error estimate (3.7) for all $p\in(2,\infty)$ and $q\in[2,\infty)$.
Step 4. Let us demonstrate the following inequality:
$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty).$ (4.57)
To this end, let $p\in(2,\infty)$ and $q\in[2,\infty)$ be arbitrarily chosen. Invoking the definitions of $II_{1}$ and $II_{2}$ from (4.55), in conjunction with Lemma 4.9 and the continuous embedding of $\dot{H}_{h}^{1,q}$ into $L^{q}$, we deduce
$\lVert{II_{1}+II_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant c\lVert{\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}.$
Relying on (4.49), (4.50), and (4.56), it follows that
$\lVert{II_{1}+II_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);L^{q})},$
which, combined with the stability property of $P_{h}$ as stated in (4.18), leads to
$\lVert{II_{1}+II_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}.$
Moreover, using Hölder’s inequality yields
$\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$
[…]
which, in light of the regularity result (4.45) and the stability assertion (4.53), simplifies to
$\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant c\lVert{y-y_{h}}\rVert_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
Consequently, we obtain the following bound for $II_{1}+II_{2}$:
$\lVert{II_{1}+II_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{y-y_{h}}\rVert_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
We proceed to bound the term $II_{3}$ from Eq. 4.55 as follows:
$\lVert{II_{3}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$ $\leqslant c\lVert{P_{h}(F(y_{h})-F(y))}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}$
$\leqslant c\lVert{F(y_{h})-F(y)}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}$
$\leqslant c\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})},$
where the first inequality is justified by Lemma 4.8 and the continuous embedding of $(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1/2-1/p,p}$ in $L^{q}$; the second inequality is a consequence of the stability of the projection $P_{h}$ from Eq. 4.18; and the final inequality relies on the Lipschitz continuity of $F$ as stated in Eq. 4.4. Combining the above bounds for $II_{1}+II_{2}$ and $II_{3}$, we arrive at
$\lVert{II_{1}+II_{2}+II_{3}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$
[…]
Given the decomposition Eq. 4.55 and the derived error estimate Eq. 3.7 for all $p\in(2,\infty)$ and $q\in[2,\infty)$, the desired inequality Eq. 4.57 is thereby established.
Step 5. We proceed to prove the error estimate Eq. 3.8. Fix arbitrarily $p\in(2,\infty)$, $q\in[2,\infty)$, and $\epsilon>0$. Using the regularity assertion Eq. 4.46 and following the proof technique used for Eq. 4.30, we deduce the following bound for $\xi_{h}$:
$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$
Employing Eq. 4.19 and the regularity assertion Eq. 4.47, we establish the following bound for $z-P_{h}z$:
$\|z-P_{h}z\|_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$
Furthermore, by combining Eqs. 4.49 and 4.57, we obtain the following bound:
$\|G-G_{h}+\eta_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}.$
Since the decomposition Eq. 4.43 implies

### v8

@@ 4.1 Preliminary Lemmas
An application of Gronwall’s inequality then yields
$\max_{t\in[0,T]}\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})}\leqslant ch^{(2-\epsilon)/p},\quad\forall\epsilon>0.$
It follows that
$\max_{t\in[0,T]}\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})}\leqslant ch^{2/p-\epsilon},\quad\forall\epsilon>0.$ (4.51)
Moreover, for any $\epsilon>0$ and $t\in(0,T]$, the following bounds hold:
$\lVert{(z-P_{h}z)(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant ch^{1-\epsilon},$ (see Eqs. 4.20 and 3.4),
@@
Now, from the decomposition Eq. 4.49, we deduce that
@@
$\qquad{}+\|\xi_{h}(t)\|_{L^{p}(\Omega;L^{2})}+\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})},\quad\forall t\in[0,T].$
Combining the above estimates, we conclude that for all $t\in(0,T]$,
$\lVert{(y-y_{h})(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant ch^{2/p-\epsilon}+ch^{1-\epsilon}+ch^{2/p}t^{-1/p},\quad\forall\epsilon>0.$
Since $h^{1-\epsilon}$ is absorbed into $h^{2/p-\epsilon}$ for $p>2$, the desired error estimate Eq. 3.6 follows immediately. This completes the proof of Theorem 3.2.
@@
Proof of Theorem 3.3. Let $G$ and $z$ be defined by Eqs. 2.8 and 4.40, respectively. Since the initial condition $y_{0}\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$ implies that $y_{0}\in\dot{H}^{2,q}$ for all $q\in(2,\infty)$, it follows from Lemma 4.2 that $G\in C([0,T];\dot{H}^{2,q})$ for all $q\in(2,\infty)$. Consequently, by adapting the argument in Step 3 of the proof of Theorem 3.1, we obtain the following improved regularity properties for the processes $y$ and $z$:
$y\in L_{\mathbb{F}}^{p}(\Omega;C([0,T];\dot{H}^{2-\epsilon,q})),$ $\forall p,q\in(2,\infty),\,\forall\epsilon\in(0,1);$ (4.52)
$z\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),$ $\forall p,q\in(2,\infty);$ (4.53)
$z\in L_{\mathbb{F}}^{p}(\Omega;C([0,T];\dot{H}^{2-\epsilon,q})),$ $\forall p,q\in(2,\infty),\,\forall\epsilon\in(0,1).$ (4.54)
Let $G_{h}$ be defined as in Eq. 4.42. We retain the notation introduced in the proof of Theorem 3.2:
@@
The remainder of the proof is structured into the following three steps.
@@
$\left\|z-P_{h}z\right\|_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p,q\in(2,\infty),$ (4.55)
$\left\|G-G_{h}\right\|_{C([0,T];L^{q})}+\left\|P_{h}G-G_{h}\right\|_{C([0,T];L^{q})}\leqslant ch^{2},\quad\forall q\in(2,\infty).$ (4.56)
Estimate (4.55) is a direct consequence of (4.20) combined with the regularity property (4.53). Estimate (4.56) follows from (4.29) and the initial condition $y_{0}\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$. Given the regularity Eq. 4.53, estimate Eq. 4.35 improves to
$\left\|\xi_{h}\right\|_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p,q\in(2,\infty),$ (4.57)
and estimate Eq. 4.36 improves to
$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-2/p-\epsilon},\quad\forall p,q\in(2,\infty),\,\forall\epsilon\in(0,1).$ (4.58)
Following the derivation of estimate Eq. 4.51 for $\eta_{h}$, we obtain the improved bound
$\max_{0\leqslant t\leqslant T}\left\|\eta_{h}(t)\right\|_{L^{p}(\Omega;L^{2})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.59)
Therefore, combining estimates Eqs. 4.57, 4.59, 4.56 and 4.55 via the decomposition Eq. 4.49 yields the error estimate
@@
Furthermore, for any $p\in(2,\infty)$, we deduce the following chain of inequalities:
$\lVert{y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{\infty})}$
@@
$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}{}$ $\lVert{P_{h}y}\rVert_{L^{p}(\Omega\times(0,T);L^{\infty})}+ch^{-3/2}\lVert{P_{h}(y-y_{h})}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}{}$ $c\lVert{y}\rVert_{L^{p}(\Omega\times(0,T);L^{\infty})}+ch^{-3/2}\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
$\stackrel{{\scriptstyle\mathrm{(iii)}}}{{\leqslant}}{}$ $c\lVert{y}\rVert_{L^{p}(\Omega\times(0,T);L^{\infty})}+ch^{1/2},$
where in (i) we apply the inverse estimate $\lVert{v_{h}}\rVert_{L^{\infty}}\leqslant ch^{-3/2}\lVert{v_{h}}\rVert_{L^{2}}$ for all $v_{h}\in X_{h}$ (see, e.g., [7, Theorem 4.5.11]); in (ii) we utilize the stability property of $P_{h}$ as stated in Eq. 4.19; and in (iii) we employ the error estimate Eq. 4.60. By Eq. 4.52 and the continuous embedding $\dot{H}^{2-\epsilon,q}\hookrightarrow L^{\infty}$ for sufficiently small $\epsilon>0$, we conclude $h$-uniform boundedness:
$\lVert{y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{\infty})}\text{ is uniformly bounded with respect to $h$ for all $p\in(2,\infty)$}.$ (4.61)
Step 2. We now prove the error estimate Eq. 3.7. The proof comprises two parts, (a) and (b).
Part (a). We establish the estimate
$\|\eta_{h}\|_{L^{p}(\Omega\times(0,T);L^{6})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.62)
Utilizing (4.44) and the definitions of the discrete convolution operators $S_{0,h}$ and $S_{1,h}$ given in Eqs. 4.30 and 4.31, we derive the decomposition
$\eta_{h}=II_{1}+II_{2},$ (4.63)
@@
$II_{1}$ $:=S_{0,h}\big{(}\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}(y^{3}-y_{h}^{3})\big{)},$
$II_{2}$ $:=S_{1,h}P_{h}\big{(}F(y_{h})-F(y)\big{)}.$
Fix $p\in(2,\infty)$. For $II_{1}$, applying Lemma 4.8(i) and Lemma 4.10 yields
$\|II_{1}\|_{L^{p}(\Omega\times(0,T);L^{\infty})}$ $\leqslant c\|II_{1}\|_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{2,5/3})}$
$\leqslant c\|\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}(y^{3}-y_{h}^{3})\|_{L^{p}(\Omega\times(0,T);L^{5/3})}$
$\leqslant c\|\xi_{h}\|_{L^{p}(\Omega\times(0,T);L^{5/3})}+c\|\eta_{h}\|_{L^{p}(\Omega\times(0,T);L^{5/3})}$
$\quad{}+c\|G_{h}-P_{h}G\|_{L^{p}(\Omega\times(0,T);L^{5/3})}+c\|P_{h}(y^{3}-y_{h}^{3})\|_{L^{p}(\Omega\times(0,T);L^{5/3})}.$
Employing the estimates Eqs. 4.59, 4.56 and 4.57 and the embedding $L^{q}\hookrightarrow L^{5/3}$ for $q>5/3$, we obtain
$\|II_{1}\|_{L^{p}(\Omega\times(0,T);L^{\infty})}\leqslant ch^{2}+\|P_{h}(y^{3}-y_{h}^{3})\|_{L^{p}(\Omega\times(0,T);L^{5/3})}.$
Since $P_{h}$ is bounded on $L^{5/3}$ (see (4.19)), it follows that
$\|II_{1}\|_{L^{p}(\Omega\times(0,T);L^{\infty})}\leqslant ch^{2}+\|y^{3}-y_{h}^{3}\|_{L^{p}(\Omega\times(0,T);L^{5/3})}.$
To estimate $\|y^{3}-y_{h}^{3}\|_{L^{p}(\Omega\times(0,T);L^{5/3})}$, we use the identity
$y^{3}-y_{h}^{3}=(y-y_{h})(y^{2}+yy_{h}+y_{h}^{2})$
and apply Hölder’s inequality, leading to
$\|y^{3}-y_{h}^{3}\|_{L^{p}(\Omega\times(0,T);L^{5/3})}\leqslant c\|y-y_{h}\|_{L^{2p}(\Omega\times(0,T);L^{2})}\Big{(}\|y\|_{L^{4p}(\Omega\times(0,T);L^{20})}^{2}+\|y_{h}\|_{L^{4p}(\Omega\times(0,T);L^{20})}^{2}\Big{)}.$
By the regularity of $y$ (see (4.52)), the stability of $y_{h}$ (see (4.61)), and the error estimate (4.60), we deduce
$\|y^{3}-y_{h}^{3}\|_{L^{p}(\Omega\times(0,T);L^{5/3})}\leqslant ch^{2}.$ (4.64)
Therefore,
$\|II_{1}\|_{L^{p}(\Omega\times(0,T);L^{\infty})}\leqslant ch^{2}.$ (4.65)
For $II_{2}$, the continuous embedding $\dot{H}^{1,2}\hookrightarrow L^{6}$ and the isometric embedding of $\dot{H}_{h}^{1,2}$ into $\dot{H}^{1,2}$ imply
$\|II_{2}\|_{L^{p}(\Omega\times(0,T);L^{6})}\leqslant c\|II_{2}\|_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{1,2})}.$
Hence, applying Lemma 4.11, the ideal property of $\gamma$-radonifying operators [16, Theorem 9.1.10], the boundedness of $P_{h}$ on $L^{2}$, the Lipschitz condition (4.4), and the estimate (4.60), we obtain
$\|II_{2}\|_{L^{p}(\Omega\times(0,T);L^{6})}$ $\leqslant c\|P_{h}\big{(}F(y_{h})-F(y)\big{)}\|_{L^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,2}))}$
$\leqslant c\|F(y_{h})-F(y)\|_{L^{p}(\Omega\times(0,T);\gamma(H,L^{2}))}$
$\leqslant c\|y-y_{h}\|_{L^{p}(\Omega\times(0,T);L^{2})}$
$\leqslant ch^{2}.$
Combining (4.65) with this estimate via the decomposition (4.63) establishes (4.62).
Part (b). We consider the decomposition of $y_{h}-y$ given in Eq. 4.49, and combine it with the estimates Eq. 4.55, Eq. 4.56, Eq. 4.57, and Eq. 4.62 to establish Eq. 3.7 for the particular case $p\in(2,\infty)$, $q=6$. Based on this, we now apply Lemma 4.8(ii) and adapt the same bounding strategy used for $II_{2}$ in part (a). This yields an improved estimate:
$\lVert{II_{2}}\rVert_{L^{p}(\Omega\times(0,T);L^{\infty})}$ $\leqslant c\lVert{II_{2}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{1,6})}$
$\leqslant c\lVert{P_{h}(F(y_{h})-F(y))}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,6}))}$
$\leqslant c\lVert{F(y_{h})-F(y)}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{6}))}$
$\leqslant c\lVert{y_{h}-y}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$
$\leqslant ch^{2},\quad\forall p\in(2,\infty).$
Combining this refined bound with the estimate for $II_{1}$ given in Eq. 4.65, and recalling the decomposition Eq. 4.63, we deduce that
@@
$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{\infty})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.66)
Finally, revisiting the decomposition Eq. 4.49 and combining the estimates Eq. 4.55, Eq. 4.56, Eq. 4.57, and Eq. 4.66, we conclude the error estimate Eq. 3.7 for all $p,q\in(2,\infty)$.
Step 3. We now establish the error estimate Eq. 3.8, proceeding in three parts, (a)–(c).
Part (a). We aim to prove the inequality
$\|\eta_{h}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}\leqslant ch^{2},\quad\forall p\in(4,\infty).$ (4.67)
To this end, fix $p\in(4,\infty)$ and let $q\in(6,\infty)$ be arbitrary. Recall the decomposition Eq. 4.63 for $\eta_{h}$. By Lemma 4.8(ii) and Lemma 4.7, we have
$\|II_{1}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant\|II_{1}\|_{L^{p}(\Omega;C([0,T];\dot{H}_{h}^{2-3/p,q}))}$
$\leqslant\|II_{1}\|_{L^{p}(\Omega;C([0,T];(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1/2-1/p,p}))}.$
An application of Lemma 4.10 then gives
$\|II_{1}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant c\|\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}(y^{3}-y_{h}^{3})\|_{L^{p}(\Omega\times(0,T);L^{q})}.$
Using the estimates Eq. 4.56, Eq. 4.57, and Eq. 4.66, we deduce
$\|II_{1}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant ch^{2}+c\|P_{h}(y^{3}-y_{h}^{3})\|_{L^{p}(\Omega\times(0,T);L^{q})}.$
Since the projection operator $P_{h}$ is bounded on $L^{q}$ (cf. Eq. 4.19), it follows that
$\|II_{1}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant ch^{2}+c\|y^{3}-y_{h}^{3}\|_{L^{p}(\Omega\times(0,T);L^{q})}.$
As in Eq. 4.64, we use the identity
$y^{3}-y_{h}^{3}=(y-y_{h})(y^{2}+yy_{h}+y_{h}^{2})$
and apply Hölder’s inequality to obtain
$\|y^{3}-y_{h}^{3}\|_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant c\|y-y_{h}\|_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
Therefore,
$\|II_{1}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant ch^{2}+c\|y-y_{h}\|_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
Next, we estimate $II_{2}$ from Eq. 4.63. Applying Lemma 4.7 and Lemma 4.8(ii), we get
$\|II_{2}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant\|II_{2}\|_{L^{p}(\Omega;C([0,T];(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1/2-1/p,p}))}$
$\leqslant c\|P_{h}(F(y_{h})-F(y))\|_{L^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,q}))},$
where the second inequality follows from Lemma 4.11. As in the estimate of $II_{2}$ in Step 2, Part (a), we then obtain
$\|II_{2}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}\leqslant c\|y-y_{h}\|_{L^{p}(\Omega\times(0,T);L^{q})}.$
Combining the bounds for $II_{1}$ and $II_{2}$, and recalling the decomposition Eq. 4.63, we arrive at
$\|\eta_{h}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant ch^{2}+c\|y-y_{h}\|_{L^{p}(\Omega\times(0,T);L^{q})}+c\|y-y_{h}\|_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
Now, since the error estimate Eq. 3.7 has already been established for all $p,q\in(2,\infty)$, we conclude that the right-hand side is of order $h^{2}$. This completes the proof of Eq. 4.67.
Part (b). We now prove the estimate
$\|y-y_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-2/p-\varepsilon},\quad\forall p,q\in(2,\infty),\ \forall\varepsilon\in(0,1).$ (4.68)
By the approximation property Eq. 4.20 and the regularity result Eq. 4.54, we have
$\|z-P_{h}z\|_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\varepsilon},\quad\forall p,q\in(2,\infty),\ \forall\varepsilon\in(0,1).$ (4.69)
Moreover, combining the approximation result Eq. 4.56 with the bound Eq. 4.67, we deduce
$\|G-G_{h}+\eta_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2},\quad\forall p,q\in(2,\infty).$ (4.70)
Using the decomposition Eq. 4.49, we write